OpenStudy (ja1):

Solve the following system by using graph paper or graphing technology. 3x – 2y = 7 9x – 6y = 3 What is the solution to the system? No Solutions Infinite Solutions (3, 2) (3, –2)

OpenStudy (anonymous):

right, if you get both of them in terms of y so 2y = 3x - 7 and 6y = 9x -3. the second can be written as 2y = 3x - 1. so then the first is y = 3/2(x) - 7/2 the second is y = 3/2(x) - 1/2 Plot those 2 lines, if they cross, the co-ordinates of the intersection is the solution. If they don't there are no solutions.

OpenStudy (anonymous):

I think that should work